### v1

@@ Abstract
The New York State Department of Transportation (NYSDOT) has a network of roadside traffic cameras that are used by both the NYSDOT and the public to observe road conditions. The NYSDOT evaluates road conditions by driving on roads and observing live cameras, tasks which are labor-intensive but necessary for making critical operational decisions during winter weather events. However, machine learning models can provide additional support for the NYSDOT by automatically classifying current road conditions across the state. In this study, convolutional neural networks and random forests are trained on camera images and weather data to predict road surface conditions. Models are trained on a hand-labeled dataset of $\sim 22{,}000$ camera images, each classified by human labelers into one of six road surface conditions: severe snow, snow, wet, dry, poor visibility, or obstructed. Model generalizability is prioritized to meet the operational needs of the NYSDOT decision makers, and the weather-related road surface condition model in this study achieves an accuracy of 81.5% on completely unseen cameras.
@@
  • Winter weather
@@
  • Artificial intelligence
  • Risk communication
  • Hand-labeled dataset
@@ 1 Introduction
Inclement weather affects traveler safety, with dangerous weather conditions leading to more than 1 million accidents in the U.S. each year (Anderson, 2024). Awareness of road surface conditions is essential for transportation agencies to know where resources (e.g., snowplows, treatments) are needed. The New York State Department of Transportation (NYSDOT) has a network of approximately 2400 live traffic cameras across New York state (NYS), which are situated at varying heights and pointing toward the roadways. These cameras (referred to as “sites” in this study) are publicly available at 511ny.org (NYSDOT, 2025) and winter road conditions are populated by the NYSDOT and displayed on a color-coded map (Fig. 1, left). Currently, the NYSDOT relies on manually observing individual cameras and driving on roads to determine the road surface conditions, both of which are time intensive. However, machine learning (ML) models, such as convolutional neural networks (CNNs) and random forests (RFs), can be used to automate the task of road surface condition classification. Compared to manual methods, the benefit of using ML is that predictions are made frequently (e.g., every five minutes) and at the individual camera level, which provides a granular spatial breakdown. When direct observation is not available, the integration of ML could enable more efficient and effective decision making because the model could automatically identify areas where roads are snow-covered and quickly accumulating (Fig. 1, right). The goal of this work is to use ML to make skillful predictions of road surface conditions on unseen camera sites (those not used in model training), prioritizing the operational needs of the NYSDOT.
@@
Weather and road surface conditions have been shown to impact driver habits (Liu et al., 2025), affect traffic and accidents (Park et al., 2024), and therefore, affect traveler safety. The Federal Highway Administration (FHA, 2023) reported that over a 10-year period (2007 to 2016), 21% percent of vehicle crashes were weather-related. Of those weather-related crashes, 70% were due to wet pavement and 16% were due to snowy/slushy pavement. With high snow rates in NYS, the impacts were even greater. From 2010 to 2019, about 35% to 50% of crashes along the NYS Thruway in the month of January were associated with inclement weather, especially snowfall (Call and Flynt, 2022). Although fog was less frequent and accounted for only 3% of weather-related crashes in the US, it accounted for 9% of weather-related fatalities (FHA, 2023), indicating the importance of recognizing poor visibility conditions. These statistics emphasize the impact of road surface conditions on driver safety, motivating the work completed in this study.
@@
Fig. 1: Existing tool and motivation. (Left) Example of the 511ny.org map on February 6, 2025. Road surface conditions are color coded according to the key provided, with most of the state showing Snow/Ice conditions (pink). (Right) Example of one camera in Buffalo, NY on November 19, 2022, with two images taken five minutes apart, demonstrating how quickly snow can accumulate on the roads.
@@ 1.1 Literature review
Research aimed at predicting road conditions from camera images has been completed using a variety of datasets and methods. Many earlier studies (Yamada et al., 2001; Kawai et al., 2014; Zhang et al., 2012; Sun and Jia, 2013; Yang et al., 2014; Lee et al., 2016; Kawarabuki and Onoguchi, 2014) employ methods of extracting features from images such as texture, edges, or hue, saturation, and value, followed by classification algorithms including K-nearest neighbors, Naïve Bayes, and Neural Networks to predict various road surface conditions, achieving accuracies ranging 87%-96%. A number of studies, including Yan et al. (2009), Omer and Fu (2010), Qian et al. (2016), Pan et al. (2018), Wu and Kwon (2022), use in-vehicle dash cameras from cars, buses, or snow plows to predict weather and road conditions. Various algorithms are used, such as Support Vector Machines and CNNs, achieving accuracies ranging 79%-93%. Other related studies incorporate information beyond road surface conditions, for example, Huang et al. (2024) use algorithms such as MonoDepth and XGBoost to predict different types of fog from roadway video data, Zhang et al. (2022) use CNNs to detect pavement distress like cracks and potholes, and Sirirattanapol et al. (2019) use multi-class CNN to predict multiple attributes including traffic levels, lighting, raining, and wet/dry road conditions. Road Weather Information System (RWIS) sensor data and camera images are used in Jonsson (2011) and Carrillo et al. (2020) to predict road surface conditions such as wet, dry, icy, and various levels of snow coverage, both works achieving accuracies above 90%. Differences in underlying datasets, classes, and applications hinder direct comparison between these past studies, however, they demonstrate that automatic road surface condition detection has been a consistent and important problem over time.
@@
Past studies that are most similar to the work presented in this paper include Chang and Walker (2022), Carrillo and Crowley (2020), and Khan and Ahmed (2022). Specifically, Chang and Walker (2022) used a dataset of 1000 NYSDOT images to predict “snow” or “no snow” conditions using Google’s Teachable Machine (Google, 2024), achieving test accuracy of 75% for the “snow” class and 95% for the “no snow” class. They also correlated image conditions with crash reports, finding that the chance of a crash increased by 61% when “snow” was predicted by the classification model. Carrillo and Crowley (2020) discussed drawbacks to using RWIS data, including their cost, which limits the number of RWIS sites, and suggests incorporation of the Ministry of Transportation of Ontario cameras as an important research consideration. Khan and Ahmed (2022) used roadside images from Wyoming DOT CCTV cameras to predict road surface conditions. Khan and Ahmed (2022) also discussed the benefits of using already-available DOT images, stating that they are more prevalent and readily available than RWIS sites given that webcams are more cost effective. They hand-labeled their image dataset with a weather condition (clear, heavy snow, and light snow) and a road surface condition (dry, snowy, and wet/slushy), and then built a separate classification model for each, achieving at least 97% validation accuracy on both. Khan and Ahmed (2022) also recognized the importance of testing their model on other roadways, so they tested their model on a small subset of 300 completely new cameras, achieving a high accuracy of 95.3%. Their work does not include poor visibility or obstruction, and the importance of these classes are included in Section 2; more comparisons to Khan and Ahmed (2022) and the work presented in this study are discussed in Section 4.1.
@@ 1.2 Contributions
This study focuses on methods that are important for building a trustworthy model (Bostrom et al., 2023) with the NYSDOT that can also be applied to broad range of scenarios and other locations. Many past studies sought to solve a similar task of automating road weather condition classification; however, some limitations of past studies include the use of one or few camera sites, the use of a limited number of road surface classes, omission of details about the underlying datasets and labeling techniques, evaluation on images from the same camera, or reliance on other equipment like in-vehicle dash cameras or expensive sensor data. Relative to past studies, five main contributions of the work presented in this paper include:
@@
    Close collaboration with the NYSDOT with the goal of creating a model that is trustworthy, a concept referred to as “co-design” (Bostrom et al., 2023).
@@ 1.3 Overview
Six road surface condition classes (severe snow, snow, wet, dry, poor visibility, and obstructed) are predicted from camera images and weather data using supervised ML models. Figure 2 shows the general flow of the modeling process: a Surface Condition Model that predicts weather-related classes using a CNN model and a downstream RF model, and an Obstruction Detection Model that uses a CNN. Section 2 discusses the datasets used in this work, including a hand-labeled image dataset and High-Resolution Rapid Refresh (HRRR) weather forecast data. Section 3 includes details on training and selecting ML models. Model performance and generalizability are discussed in Section 4. The terms “prediction” and “predict” are used to refer to a model’s classification output. A primary focus of this work is to understand how the model performs on unseen camera sites, referred to as “generalizability”. Generalizability is critical for meeting operational needs that involve deploying the road surface classification models on thousands of cameras across NYS.
Fig. 2: Operational model flow and example. (Left) The Surface Condition Model is a two-stage process using a CNN and a RF. The Obstruction Detection Model solely predicts whether an image is obstructed of non-obstructed. (Right) The model flow in a practical inference scenario is shown, demonstrating that both predictions will be provided in conjunction for each new observation.
@@ 2.1 Image and Weather Data
Camera images are utilized to create a labeled dataset, which serves as the foundation for developing ML models in this study. Images are viewed manually by human labelers who assign a road surface condition classification to each image. A systematic approach from social science, known as Quantitaive Content Analysis (Coe and Scacco, 2017), is used to label the images in this study. This approach includes the development of a codebook, which serves as thorough documentation of classification rules and definitions to ensure replicability and reproducibility, as published in Wirz et al. (2024). The New York State Mesonet (NYSM; Brotzge et al., 2020) observational weather data are used to aid in hand-labeling. Comprehensive documentation for the data used in this study is available at Sutter et al. (2025b); a total of 21,653 images from 45 unique camera sites across NYS are labeled for this study. Table 1 shows examples of the six labeled classes, and detailed definitions for each class are provided in the codebook from Sutter et al. (2025a).
@@
A large focus of this work is to create a model that performs well on a wide range of conditions that are realistically encountered in the installed cameras. For example, the obstructed cases seem less important because they are generally not weather related, but nontheless, they are included because they are observed frequently and can deem the observation as unreliable. Similarly, cameras that are of lower quality but are not obstructed (right-hand side of Table 1) are included in this study. While past studies may have included difficult examples like these, they have not been emphasized, and the focus of this work is to provide a model that can work on representative images from already-implemented cameras, which include these less-than-ideal cases. Building one model that can predict road surface conditions across a wide range of conditions on all cameras (including low quality cameras), creates a tool that is operationally realistic for the NYSDOT.
@@
Table 1: The six road surface conditions are shown with examples and count distribution, totaling to 21,653 images. The right-most column shows examples of high- and low-quality images. The high-quality image is well lit even in the absence of car headlights, making the classification of road surface conditions more obvious. The low-quality image is poorly lit and is distanced far from the roadway, making the classification of these images more difficult; in these cases, the labeler relies on slight changes in road pavement darkness as well as NYSM observational data (Sutter et al., 2025a).
. Six classes Site Quality Severe snow - 1,527 Snow - 2,989 Wet - 8,458 High Quality - 58% [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Dry - 7,600 Poor visibility - 915 Obstructed - 164 Low Quality - 42% [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
The High-Resolution Rapid Refresh (HRRR, 2024) is a weather forecast model providing information like forecasted snowfall, rainfall, temperature, and wind speeds. HRRR forecast data are updated hourly and used in a wide range of applications where weather predictions are needed (Alexander et al., 2013). The labeled image data and forecast data are collated based on time and location resulting in corresponding forecast data for each labeled image. The forecast data being used is forecast hour two, with valid time corresponding to the time of the paired labeled image. The HRRR data are used in favor of observational data because of its fine spatial grid of 3km. The average distance from each camera to the nearest HRRR grid point is 1.1km. Six HRRR weather forecast variables are used in this study, including 2-meter temperature, 2-meter relative humidity, 10-meter average wind speed, 2-hr accumulated snow depth, 2-hr total precipitation, and total cloud cover, but other HRRR data variables should be considered in future work. Since five of the labeled classes in this study are weather-related, it is reasonable to include weather data in the modeling process for the Surface Condition Model, and doing so made a positive impact to model performance, as discussed in Section 4.3. The obstructed class is generally not weather-related, so HRRR data are not used for the Obstruction Detection Model (see Fig. 2).
@@ 2.2 Data splitting
The data is split to ensure that model evaluation aligns with the goal of generalizability. Specifically, a “site-specific” split method is used where all observations from a site are in the same fold, which ensures that training, validation, and testing contain unique sites.
Nested cross-validation (CV) is used for the Surface Condition Model to ensure that each stage of the modeling process is evaluated on data not used in model training. The labeled data is split into six folds of approximately equal size (16.67%) that serve as the training data, validation data, or testing data at different points in the modeling process. Training dataset 1 is used to train the CNN, and training dataset 2 is used to train the RF model; the benefit of using separate datasets are discussed in Sections 3.2.2 and 4.3. Additionally, the nested structure allows for a more robust evaluation of ensemble methods (compared to a single test set) because each fold serves as the test dataset (see Section 3.4). Data splitting is demonstrated visually in Fig. 3.
Fig. 3: Nested CV for the Surface Condition Model. For each test dataset (black cells), inner 5-fold CV is used on the remaining five folds (right), resulting in five models. The five models are used to create an ensemble prediction for the test dataset. This process is demonstrated for Test 1 and is repeated for each test set (left).
For the Obstruction Detection Model, there are only 164 examples of obstructed images. Three sets of non-obstructed images are sampled from the 5-class dataset, and undersampling is done purposefully to be approximately equal to the number of obstructed samples, while also representing observations from each class and site from the 5-class dataset, resulting in 194 non-obstructed images. In total, the two-class dataset for the Obstruction Detection Model contains only 358 images, and two-fold CV is used due to the limited number of observations. Two-fold CV provides the training and validation datasets only. Generally, it is not a best practice to omit a testing dataset, however, for the Obstruction Detection Model, the validation data is only used so much as to stop model training before overfitting occurs; no model testing or hyperparameter tuning is done (see Sections 3.2.3 and 3.3). Thus, using two-fold CV provides a practical solution given the limited size of the dataset. The purpose of three different samples of non-obstructed images is to provide a framework for ensembling. Although the obstructed examples are the same between each of the three datasets, having different non-obstructed examples makes for three different datasets that are trained separately and used to create a 3-member ensemble. This approach provides a way of doing ensembling without nested CV, which is impractical with this small dataset. Data splitting for the Obstruction Detection is shown in Fig. 4.

### v2

@@ Abstract
Transportation agencies make critical operational decisions during hazardous weather events, including assessment of road conditions and resource allocation. In this study, machine learning models are developed to provide additional support for the New York State Department of Transportation (NYSDOT) by automatically classifying current road conditions across the state. Convolutional neural networks and random forests are trained on NYSDOT roadside camera images and weather data to predict road surface conditions. This task draws critically on a robust hand-labeled dataset of $\sim 22{,}000$ camera images containing six road surface conditions: severe snow, snow, wet, dry, poor visibility, or obstructed. Model generalizability is prioritized to meet the operational needs of the NYSDOT decision makers, including integration of operational datasets and use of representative and realistic images. The weather-related road surface condition model in this study achieves an accuracy of 81.5% on completely unseen cameras. With operational deployment, this model has the potential to improve spatial and temporal awareness of road surface conditions, which can strengthen decision-making for operations, roadway maintenance, and traveler safety, particularly during winter weather events.
@@
  • Machine learning
  • Traffic camera imagery
  • Winter road maintenance
@@
  • New York State Department of Transportation
@@ 1 Introduction
Intelligent systems and use of data-driven approaches have been critical for many parts of the transportation sector, including traffic safety (diverging), traffic flow prediction (trafficflow), construction of new roadways, preventative maintenance of bridges, traffic management, urban mobility and sustainability, and public transportation (OlivieraAvenues; OliveiraBridges; OlivieraUrban; OliveiraPublicTransportation; OlivieraTraffic). Technology has played an essential role in developing unique solutions for transportation, and this work focuses on detecting road surface conditions that could support operational decision-making and maintenance, especially during winter months.
@@
Inclement weather affects traveler safety, with dangerous weather conditions leading to more than 1 million accidents in the U.S. each year (ncar-road-conditions). Weather and road surface conditions have been shown to impact driver habits (Liu), affect traffic and accidents (Park), and therefore, affect traveler safety. The Federal Highway Administration (fha) reported that over a 10-year period (2007 to 2016), 21% percent of vehicle crashes were weather-related. With high snow rates in New York State (NYS), the impacts were even greater. From 2010 to 2019, about 35% to 50% of crashes along the NYS Thruway in the month of January were associated with inclement weather, especially snowfall (CallFlynt). Although fog was less frequent and accounted for only 3% of weather-related crashes in the US, it accounted for 9% of weather-related fatalities (fha). These statistics indicate the impact of road surface conditions on driver safety, motivating the work completed in this study.
Awareness of road surface conditions is essential for transportation agencies to know where resources (e.g., snowplows, treatments) are needed. The New York State Department of Transportation (NYSDOT) has a network of approximately 2400 live traffic cameras across NYS, which are situated at varying heights and pointing toward the roadways. These cameras (referred to as “sites” in this study) are publicly available at 511ny.org (ny511) and winter road conditions are populated by the NYSDOT. Currently, the NYSDOT relies on manually observing individual cameras and driving on roads to determine the road surface conditions, both of which are time-intensive given the extensive road network across NYS (Fig. 1, left). However, machine learning (ML) models, such as convolutional neural networks (CNNs) and random forests (RFs), can be used to automate the task of road surface condition classification. Compared to manual methods, the benefit of using ML is that predictions are made frequently (e.g., every five minutes) and at the individual camera level, which provides a granular spatial breakdown. When direct observation is not available, the integration of ML could enable more efficient and effective decision making because the model could automatically identify areas where roads are snow-covered and quickly accumulating (Fig. 1, right). The purpose of this work is to use ML to make skillful predictions of road surface conditions on unseen camera sites (those not used in model training), prioritizing the operational needs of the NYSDOT.
@@
Fig. 1: Roadways and motivation. (Left) Topological road network graph of New York State. (Right) Example of one camera in Buffalo, NY on November 19, 2022, with two images taken five minutes apart, demonstrating how quickly snow can accumulate on roads.
@@ 1.1 Literature review
Camera images are used in transportation research across a variety of applications, using a range of datasets and methods. Many earlier studies (Yamada2005; Kawai2014; ZhangChunyu2012; Sun2013; hunjunyang2014; Lee2016; kawarabuki2014) employ methods of extracting features from images such as texture, edges, or hue, saturation, and value, followed by classification algorithms including K-nearest neighbors, Naïve Bayes, and Neural Networks to predict various road surface conditions, achieving accuracies ranging 87%-96%. A number of studies, including Yan2009, omer2010, qian2016, pan2018, wu2022, use in-vehicle dash cameras from cars, buses, or snow plows to predict weather and road conditions. Various algorithms are used, such as Support Vector Machines and CNNs, achieving accuracies ranging 79%-93%. Other related studies incorporate information beyond road surface conditions, for example, Huang use algorithms such as MonoDepth and XGBoost to predict different types of fog from roadway video data, ZHANG2022298 use CNNs to detect pavement distress like cracks and potholes, ports develop a novel algorithm for autonomous navigation in ports, and sirirattanapol use multi-class CNN to predict multiple attributes including traffic levels, lighting, raining, and wet/dry road conditions. Road Weather Information System (RWIS) sensor data and camera images are used in jonsson and carrillo2020design to predict road surface conditions such as wet, dry, icy, and various levels of snow coverage, both works achieving accuracies above 90%. Differences in underlying datasets, classes, and applications hinder direct comparison between these past studies, however, they demonstrate that automatic road surface condition detection has been a consistent and important problem over time.
@@
Past studies that are most similar to the work presented in this paper include walker2022, carrillo2020integration, and KHAN2022. Specifically, walker2022 used a dataset of 1000 NYSDOT images to predict “snow” or “no snow” conditions using Google’s Teachable Machine (teachablemachine), achieving test accuracy of 75% for the “snow” class and 95% for the “no snow” class. They also correlated image conditions with crash reports, finding that the chance of a crash increased by 61% when “snow” was predicted by the classification model. carrillo2020integration discussed drawbacks to using RWIS data, including their cost, which limits the number of RWIS sites, and suggests incorporation of the Ministry of Transportation of Ontario cameras as an important research consideration. KHAN2022 used roadside images from Wyoming DOT CCTV cameras to predict road surface conditions. KHAN2022 also discussed the benefits of using already-available DOT images, stating that they are more prevalent and readily available than RWIS sites given that webcams are more cost effective. They hand-labeled their image dataset with a weather condition (clear, heavy snow, and light snow) and a road surface condition (dry, snowy, and wet/slushy), and then built a separate classification model for each, achieving at least 97% validation accuracy on both. KHAN2022 also recognized the importance of testing their model on other roadways, so they tested their model on a small subset of 300 completely new cameras, achieving a high accuracy of 95.3%. Their work does not include poor visibility or obstruction, and the importance of these classes are included in Section 2; more comparisons to KHAN2022 and the work presented in this study are discussed in Section 4.1.
@@ 1.2 Contributions
This study focuses on methods for building a trustworthy model (bostrom2023trust) with NYSDOT that can generalize across scenarios and locations. While many past studies addressed automated road-weather condition classification, they were often limited by few camera sites, few surface classes, limited dataset and labeling details, same-camera evaluation, or reliance on in-vehicle cameras or expensive sensors. Relative to past studies, this work offers five key contributions:
@@
    Close collaboration with the NYSDOT with the goal of creating a model that is trustworthy, a concept referred to as “co-design” (bostrom2023trust).
@@ 1.3 Overview
In this work, six road surface condition classes (severe snow, snow, wet, dry, poor visibility, and obstructed) are detected from camera images and weather data using supervised ML models. As shown in Figure 2, the workflow consists of a Surface Condition Model, which predicts weather-related classes using a CNN followed by an RF, and an Obstruction Detection Model using a CNN. Section 2 describes the hand-labeled image dataset and HRRR weather forecast data, and Section 3 outlines model training and selection. Section 4 evaluates model performance, with particular attention to performance on unseen camera sites, referred to as “generalizability”. Generalizability is critical for meeting operational needs that involve deploying the road surface classification models on thousands of cameras across NYS.
Fig. 2: Operational model flow and example. The Surface Condition Model is a two-stage process using a CNN and a downstream model (RF). The Obstruction Detection Model predicts whether an image is obstructed or non-obstructed. In a practical inference scenario, both predictions will be provided in conjunction for each new observation.
@@ 2.1 Image and Weather Data
A labeled dataset has been developed for this study for the development of ML models. This dataset consists of hand-labeled classifications of road surface conditions, which came from human labelers who manually viewed camera images and assigned a classification to each. Following a systematic approach known as Quantitative Content Analysis (qcaorig), this process included the development of a codebook containing documentation and rules for labeling each class (Sutterzenodo), where definitions were derived in consultation with the NYSDOT; for example, the presence of tire tracks distinguish between snow and severe snow. Critically, this labeling process also included a set of labeling trials where six people independently labeled the same set of images to ensure reliability, and replicability of this hand-labeled dataset. Comprehensive details regarding the dataset for this study are published and available at Wirz2024, Sutterzenodo, and Sutterdatasheet. A total of 21,653 images from 45 unique camera sites across NYS are labeled for this study. Table 1 shows examples of the six labeled classes, and detailed labeling definitions are in the codebook (Sutterzenodo).
@@
A large focus of this work is to create a model that performs well on a wide range of conditions that are realistically encountered in the installed cameras, such as lower quality cameras (right-hand side of Table 1). Additionally, the obstructed cases seem less important because they are generally not weather related, but are included because they are observed frequently and can deem the observation as unreliable. While past studies may have included difficult examples like these, they have not been emphasized, and the focus of this work is to provide a model that can work on representative images from already-implemented cameras, which include these less-than-ideal cases. Building one model that can predict road surface conditions across a wide range of conditions on all cameras (including low quality cameras), creates a tool that is operationally realistic for the NYSDOT.
@@
Table 1: The six road surface conditions are shown with examples and count distribution, totaling to 21,653 images. The right-most column shows examples of high- and low- quality images. The high-quality image is well lit even in the absence of car headlights, making the classification of road surface conditions more obvious. The low-quality image is poorly lit and is distanced far from the roadway, making the classification of these images more difficult; in these cases, the labeler relies on slight changes in road pavement darkness an observational weather data (Sutterzenodo).
Six classes Site Quality
Severe snow - 1,527 Snow - 2,989 Wet - 8,458 High Quality - 58%
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Dry - 7,600 Poor visibility - 915 Obstructed - 164 Low Quality - 42%
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
The High-Resolution Rapid Refresh (HRRR) (hrrr) is a weather forecast model specifically designed to meet the demands of applications that depend heavily on timely and accurate short-range meteorological data (hrrr2013). The labeled image data and forecast data are collated based on time and location resulting in corresponding weather data for each labeled image. Although forecasts contain errors compared to observational weather stations, such as the New York State Mesonet (NYSM) (nysm), forecast data is selected for this study for multiple reasons. First, the HRRR has a fine spatial grid with an average of 3km between HRRR grid points, whereas for the NYSM data, the average spacing is 27km. Second, with the broad goal to expand to other states and regions, the HRRR has large spatial coverage available nationwide, whereas relying on states to have expensive weather-observation stations could limit expansion and applicability. Lastly, the HRRR model uses frequent data assimilation to incorporate recent observations such as radar, producing highly reliable short-term forecasts (gaudet), and this study capitalizes on that reliability by utilizing a short lead time of 2 hours while also avoiding complications associated with initial spinup. Overall, weather data makes a positive contribution to the Surface Condition Model, resulting in an overall accuracy improvement of +12.1% (Sections 4.1 and 4.3).
Six near-surface HRRR variables are used in this study including 2-meter temperature, 2-meter relative humidity, 10-meter average wind speed, 2-hr accumulated snow depth, 2-hr total precipitation, and total cloud cover. These six variables are selected based on their prominent relevance to road-surface weather and are similar to those used in carrillo2020design. The HRRR data are not used to predict obstructions given that the obstructed class is generally not weather-related.
@@ 2.2 Data splitting
The data is split to ensure that model evaluation aligns with the goal of generalizability; this is critical given that labeling images from every camera site is not practical. Specifically, a “site-specific” split method is used where all observations from a site are in the same fold, which ensures that training, validation, and testing contain unique sites. Folds are created to optimize the distribution of multiple attributes, although exactly equal composition is not feasible given the iterative nature of labeling (Sutterdatasheet). Specifically, each of the six folds are comprised of between 1) 22-29% hazardous classes (severe snow, snow, or poor visibility), 2) 6-8 unique sites, 3) 38-55% low quality sites 4) 70-90% from the months of January or February, and 5) 52-56% night instances. Winter months are prioritized for operational purposes, but the same model works in summer months, predicting wet, dry, and poor visibility classes. Data splitting for the Surface Condition Model is demonstrated visually in Fig. 3, showing how nested cross-validation (CV) is used to ensure that each stage of the modeling process is evaluated on data not used in model training. The labeled data is split into six folds of approximately equal size (16.67%) that serve as the training, validation, or testing data at different points in the modeling process. Training dataset 1 is used to train the CNN, and training dataset 2 is used to train the stage 2 downstream model; the benefits of using separate datasets are discussed in Sections 3.2.2 and 4.3. The nested structure allows for a robust evaluation of ensemble methods compared to a single test set (see Section 3.4).
Fig. 3: Nested CV for the Surface Condition Model. For each test dataset (black cells shown on the left), inner 5-fold CV is used on the remaining five folds (right), resulting in five models. The five models are used to create an ensemble prediction for the test dataset. This process is demonstrated for Test 1 and is repeated for each test set.
For the Obstruction Detection Model, only 164 obstructed images are available. As shown in Fig. 4, three sets of non-obstructed images are sampled from the 5-class dataset to roughly match the non-obstructed count while still representing all classes and sites, yielding 194 non-obstructed images. Due to the small dataset containing 358 samples total, two-fold CV is used to provide training and validation splits only. Although omitting a test set is generally not a best practice, the validation data is used solely for early stopping; no model testing or hyperparameter tuning is performed (see Sections 3.2.3 and 3.3), making two-fold CV a practical choice for this small dataset. The three non-obstructed samples enable a 3-member ensemble: the obstructed images remain the same across datasets, but the varied non-obstructed samples produce three different datasets trained separately for the 3-member ensemble. This allows ensembling without nested CV, which is impractical for such a small dataset.